OpenStudy (anonymous):

Solve the equation for x: 2x - 16y = 10. A. x = 10 + 8y B. x = 5 - 16y C. x = 5 + 8y D. x = 5 + 16y Solve the equation for x: 2x - 16y = 10. A. x = 10 + 8y B. x = 5 - 16y C. x = 5 + 8y D. x = 5 + 16y @Mathematics

OpenStudy (anonymous):

ok this is how you do this 2x-16y=10 +16y on both side to isolate the x 2x=10 + 16y now divide everything by 2 isolate the x to just have x x=(10+16y)/2 this also means x=(10)/2 + 16y/2 which can be simplified to x=5 + 8y therefore C

OpenStudy (anonymous):

does that help??

OpenStudy (anonymous):

Yes it does thank you very much for your help

OpenStudy (anonymous):

you are very welcome, if you have any questions feel free to message me